OpenStudy (anonymous):

Find the sum of the first 12 terms of the sequence. Show all work for full credit. 1, -4, -9, -14, . . . please this is my last question and i cant seem to get it right. if possible can someone work it out for me?

OpenStudy (campbell_st):

what do you know about the question..?

OpenStudy (anonymous):

i believe its an arithmetic sequence?

OpenStudy (campbell_st):

good start, what's the common difference..?

OpenStudy (anonymous):

ok that where i confuse myself

OpenStudy (campbell_st):

well what is the sequence changing by... common difference = 2nd term - 1st term check with 3rd term - 2nd term.... is it the same

OpenStudy (anonymous):

-5

OpenStudy (campbell_st):

great... and the 1st term a = 1 so for the sum of the sequence you can use the formula \[S_{n} = \frac{n}{2}[2a + (n - 1) \times d]\] you are given a = 1, you found d = -5 and have been told n = 12 so substitute them and calculate the sum

OpenStudy (anonymous):

i get the plugging in but what does d equal? all the rest i plugged in

OpenStudy (campbell_st):

d is the common difference d = -5

OpenStudy (anonymous):

i see ok 1 min

OpenStudy (anonymous):

is it 6/-53

OpenStudy (anonymous):

that's what i got but I not entirely sure.

OpenStudy (campbell_st):

you are close the calculation becomes \[S_{12} = \frac{12}{2} \times (-53)\]

OpenStudy (campbell_st):

so then \[S_{12} = 6 \times (-53)\]

OpenStudy (anonymous):

so that is the final form?